### v1

# Emergent SU(3) topological system in a trimer SSH model

Abstract
We consider a trimer Su-Schrieffer-Heeger (SSH) tight-binding Hamiltonian keeping up to next-nearest-neighbor (NNN) hopping terms and on-site potential energy. The Bloch Hamiltonian can be expressed in terms of all the eight generators (i.e. Gell-Mann matrices) of the SU(3) group. We provide exact analytical expressions of three dispersive energy bands and the corresponding eigenstates for any choices of the system parameters. The system lacks full chiral symmetry since the energy spectrum is not symmetric around zero, except at isolated Bloch wavevectors. We explore parity, time reversal, and certain special chiral symmetries for various system parameters. We discuss the bulk-boundary correspondence by numerically computing the Zak phase for all the bands and the boundary modes in the open boundary condition. There are three different kinds of topological phase transitions which are classified based on the gap closing points in the Brillouin zone (BZ) while tuning the nearest-neighbor (NN) and NNN hopping terms. We find that quantized changes (in units of $\pi$) in two out of three Zak phases characterize these topological phase transitions. We propose another bulk topological invariant, namely the sub-lattice winding number, which also characterizes the topological phase transitions changing from $\nu^{\alpha}=0\leftrightarrow 2$ and $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ ($\alpha$: sub-lattice index). The sub-lattice winding number provides a relatively simple analytical understanding of topological phases and may help in characterizing topological phases of systems without chiral symmetry.
I Introduction
The geometrical and topological aspects of the Bloch bands in solid state systems have attracted intense research activities [1, 2, 3, 4, 5, 6]. The symmetry-protected topological (SPT) phases of matter can be understood in terms of the interplay between symmetry and topology [7, 8]. For non-interacting systems, the SPT phases are classified in the 10-fold periodic table according to the system’s spatial dimension and internal symmetries which include time reversal, particle-hole, and chiral symmetry [9]. The bulk-boundary correspondence is the underlying principle of topological phases which relates the non-trivial bulk topological invariant to the number of topological boundary states in the open boundary condition. Most of the studies are carried out in two-dimensional systems (2D) [10, 11], however the studies have been extended to one-dimensional (1D) [12, 13] and three-dimensional (3D) systems [14] as well.
The simplest model describing the topological features is the dimer Su-Schrieffer-Heeger (SSH) model [15] consists of a 1D tight-binding model with alternating hopping parameters. Further, topological phases of the extended SSH model including the next-nearest-neighbour (NNN) hopping has been studied extensively [16]. The topological phase of a finite system can be characterized by the presence of zero energy boundary modes which are exponentially localized at the edges. The conventional bulk-boundary correspondence using the Zak phase [17], similar to the Berry phase [18], has been used as a topological number to classify various inversion-symmetric 1D topological systems where it is quantized to $\pi$ or 0 (mod $2\pi$) [19]. The boundary modes and the Zak phase in various artificial 1D systems such as optical lattices [20], acoustic crystals [21], photonic lattices [22], photonic crystals [23, 24] and artificial spin chains in ultracold atomic system [25] have been studied elaborately. However, there is an ambiguity to use the Zak phase as a topological number, since it depends on the choice of the origin in the real-space and how the unit cell is chosen [20, 26]. For systems in odd spatial dimensions with chiral and/or inversion symmetry, three bulk-topological invariants namely the winding number, and inversion symmetry indicator are proposed [27].
The momentum space Hamiltonian of 1D, 2D and 3D topological systems having two bands is represented by irreducible SU(2) generators. However, there is a growing interest in three-band in 2D and 3D topological systems, whose momentum space Hamiltonian is represented by an irreducible SU(3) generators [28, 29] given in Appendix A. For example, Lieb [30, 31, 32] and $\alpha$-$\mathcal{T}_{3}$ lattices [33, 34] in 2D as well as triple-component fermions [35, 36, 37, 38] in 3D exhibits three bands, two dispersive bands, and one flat band. All three bands touch at a given ${\bm{k}}$ point in the Brillouin zone (BZ), leading to a three-fold degeneracy at the band touching point. In Ref. [39], it was proposed that in a square optical lattice with spin-orbit coupling subjected to a spatially homogeneous SU(3) synthetic gauge field can exhibit nontrivial topology. The possibility of designing a 2D lattice model having SU(3) symmetry for spinless fermion has been discussed in Ref. [40]. In a recent study [41], a two-dimensional electron gas with Rashba spin-orbit coupling has been generalized to an SU(3) system, by considering trigonal crystal field effect which leads to a partial splitting of the energy levels at the time-reversal invariant momenta. But there are very few studies [42, 43] on three-band topological systems in 1D.
In this work, we consider a trimer SSH model by considering three sites in a given unit cell. Keeping on-site energy and hopping term up to NNN sites, the Bloch Hamiltonian is expressed in terms of all the eight generators of the SU(3) group. We obtain exact analytical expressions of three dispersive energy bands and the associated eigenstates for arbitrary choices of the system parameters. We explore various symmetry properties of the trimer SSH model for various system parameters. Since the system does not possess chiral symmetry, we establish the bulk-boundary condition by numerically computing the Zak phases ($\phi_{\lambda}$ with $\lambda=1,2,3$) for all the bands along with the boundary modes in the open boundary condition. Depending on the gap closing points in the BZ, we observe three different kinds of topological phase transitions while tuning NN and NNN hopping terms. We find that quantized changes in two out of three Zak phases are necessary to characterize these topological phase transitions, namely $\phi_{1}+\phi_{3}$ changing from $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ as well as $\phi_{1}+\phi_{3}=0\leftrightarrow\pi\leftrightarrow 2\pi$ characterize these topological phase transitions. Moreover, we propose another topological invariant, namely the sub-lattice winding number which also characterizes topological phase transitions similar to the Zak phases. We derive exact analytical expressions of the sub-lattice winding number which changes from $\nu^{\alpha}=0\leftrightarrow 2$ as well as $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ across the topological phase transitions.
This paper is organized as follows. In Sec. II, we describe the trimer SSH model and obtain the periodic Bloch Hamiltonian, dispersive energy bands along with the corresponding eigenstates and various symmetries of the system. In Sec. III, we discuss Zak phases, sub-lattice winding number, and the boundary modes in the open boundary condition. We summarize our results in Sec. IV. In Appendix A, we provide an explicit form of the Gell-Mann matrices. We obtain canonical Bloch Hamiltonian and show the equivalence with the periodic Bloch Hamiltonian in Appendix B. The phase diagram characterizing topological phases in different parameter regimes is given in Appendix C. The sub-lattice winding number for the dimer SSH model is discussed in Appendix D.
II SU(3) Hamiltonian and its various properties
We consider a toy model of a 1D chain with three sites, labelled by $A,B$ and $C$, in each unit cell as shown in Fig. 1. Considering nearest-neighbor (NN) and NNN hopping terms along with the on-site interaction energies ±$\pm\Delta$ at the sites $A$ and $C$ respectively, the trimer SSH tight-binding Hamiltonian in real space reads
$H$ $=$ $\sum_{l}[ua_{l}^{\dagger}b_{l}+vb_{l}^{\dagger}c_{l}+wa_{l+1}^{\dagger}c_{l}+h.c.]$ (1)
$+$ $\sum_{l}[w_{1}a_{l}^{\dagger}c_{l}+u_{1}b_{l}^{\dagger}a_{l+1}+v_{1}c_{l}^{\dagger}b_{l+1}+h.c.]$
$+$ $\Delta\sum_{l}[a_{l}^{\dagger}a_{l}-c_{l}^{\dagger}c_{l}].$
Here the operators $a_{l}^{\dagger}$, $b_{l}^{\dagger}$ and $c_{l}^{\dagger}$ create a spinless fermion at the sub-lattice sites $A,B$ and $C$ in the $l$-th unit cell of unit length and $h.c.$ stands for the hermitian conjugate. As opposed to the dimer SSH model, the NNN hopping occurs between two different sites in the same cell as well as in the next cell. In the absence of NNN hopping amplitudes and on-site energy, the above Hamiltonian $H$ reduces to that of in Ref. [42, 43].
Figure 1: Schematic diagram of a trimer SSH model. The unit cell of unit length is encircled by the dotted rectangle containing three sites $A,B$, and $C$. Here, $u$, $v$ are the intra-cell NN hopping terms and $w$ is the inter-cell hopping term. Similarly, $u_{1}$, $v_{1}$ are the inter-cell NNN hopping terms and $w$ is the intra-cell NNN hopping term.
We will discuss the Bloch Hamiltonian of the trimer SSH model using two different basis states, a periodic Bloch Hamiltonian and the canonical Bloch Hamiltonian or simply Bloch Hamiltonian which will be discussed in Appendix B. Both representations may be useful and convenient to study certain aspects of the systems.
Periodic Bloch Hamiltonian: Due to translational invariance, the Hamiltonian can be diagonalized using the cell index $l$. The real space field operators at the sub-lattice sites $A,B$ and $C$ in the $l$-th cell can be expanded as
$(a_{l},b_{l},c_{l})$ $=$ $\frac{1}{\sqrt{N}}\sum_{k}e^{ilk}(a_{k},b_{k},c_{k}),$ (2)
where $k$ is the Bloch wave vector restricted within the first BZ between $-\pi$ to $\pi$ and $N$ is the number of unit cell. There are $3N$ lattice sites in the whole chain. Using the properties $\sum_{l}e^{il(k-k^{\prime})}=N\delta_{k,k^{\prime}},$ the Hamiltonian reduces to the form
$H$ $=$ $\sum_{k}\psi_{k}^{\dagger}\mathcal{H}(k)\psi_{k},$ (3)
where the spinor $\psi_{k}^{\dagger}=(a_{k}^{\dagger},b_{k}^{\dagger},c_{k}^{\dagger})$ and the periodic Bloch Hamiltonian $\mathcal{H}(k)$ is given by
$\mathcal{H}(k)$ $=$ $\left(\begin{array}[]{ccc}\Delta&U&W\\ U^{*}&0&V\\ W^{*}&V^{*}&-\Delta\end{array}\right),$ (7)
where the complex elements are $U=u+u_{1}e^{-ik}$, $V=v+v_{1}e^{-ik}$ and $W=w_{1}+we^{-ik}$. In real space, the trimer SSH Hamiltonian is a $3N\times 3N$ matrix, whereas it becomes a $3\times 3$ matrix $\mathcal{H}(k)$ for each value of $k$. While performing the Fourier transformation of the Hamiltonian, only the cell number $l$ is used in the Fourier summation (see Eq. 2), without mentioning the actual position of the sublattices within each cell. As a result, this Hamiltonian is periodic in $2\pi$: $\mathcal{H}(k+2\pi)=\mathcal{H}(k)$.
The Hamiltonian $\mathcal{H}(k)$ can be expressed in terms of all the SU(3) generators, eight Gell-Mann matrices given in Appendix A, as
$\mathcal{H}(k)$ $=$ ${\bm{\Lambda}}\cdot{\bm{d}(k)}.$ (8)
Here ${\bm{\Lambda}}$ is a vector composed of the Gell-Mann matrices and ${\bm{d}(k)}$ is an eight-dimensional real vector arising due to various hopping terms up to NNN sites in 1D, whose components are $d_{1}=u+u_{1}\cos(k),d_{2}=u_{1}\sin(k),d_{3}=\Delta/2,d_{4}=w_{1}+w\cos(k),d_{5}=w\sin(k),d_{6}=v+v_{1}\cos(k),d_{7}=v_{1}\sin(k)$ and $d_{8}=\sqrt{3}\Delta/2$. The length of the eight-dimensional vector ${\bm{d}(k)}$ is $L(k)=\sqrt{\Delta^{2}+|U|^{2}+|V|^{2}+|W|^{2}}$. In the absence of the NNN hopping, $L$ becomes $k$-independent, which is in sharp contrast to the dimer SSH model.
II.1 Energy bands and their properties
The Hamiltonian, $\mathcal{H}(k)$ yield the following depressed cubic equation for eigenvalues $E$: $E^{3}+pE+q=0$ with
$p$ $=$ $-[(\Delta^{2}+u^{2}+v^{2}+w^{2}+u_{1}^{2}+v_{1}^{2}+w_{1}^{2})+2(uu_{1}+vv_{1}+ww_{1})\cos(k)]=-L^{2}(k)$ (9)
$q$ $=$ $-[\Delta(u^{2}-v^{2}+u_{1}^{2}-v_{1}^{2})+2(u_{1}vw+uv_{1}w+uvw_{1})+2u_{1}v_{1}w_{1}\cos(2k)$ (10)
$+$ $2(\Delta(uu_{1}-vv_{1})+uvw+u_{1}v_{1}w+u_{1}vw_{1}+uv_{1}w_{1})\cos(k)].$
It should be noted that $p$ is always negative since $p=-L^{2}(k)$. The three energy bands are
$E_{\lambda}(k)=2\sqrt{\frac{-p}{3}}\cos\left[\frac{1}{3}\arccos\left(\frac{3q}{2p}\sqrt{\frac{-3}{p}}\right)+\frac{2\pi\lambda}{3}\right]$ (11)
with $\lambda=1,2,3$ and $E_{3}\geq E_{2}\geq E_{1}$. It is important to note here that the middle band $E_{2}(k)$ vanishes at two $k$ points whenever $q=0$. Therefore the two locations of the zeros of $E_{2}(k)$ can be found from Eq. (10) for the given system parameters. For example, $E_{2}(k=\pm\pi/2)=0$ for the following two simple cases: (i) $\Delta\neq 0$, $u=v$ and $u_{1}=v_{1}=w_{1}=0$ and (ii) for any values of $u,v,w$ with $\Delta=u_{1}=v_{1}=w_{1}=0$. Figures 2 and 3 display the band dispersion for the various system parameters.
The corresponding eigenspinors are obtained as
$\psi_{\lambda}(k)=N_{\lambda}\left(\begin{array}[]{c}(E_{\lambda}+\Delta)U+V^{*}W\\ E_{\lambda}^{2}-\Delta^{2}-|W|^{2}\\ (E_{\lambda}-\Delta)V^{*}+UW^{*}\end{array}\right),$ (12)
where the $N_{\lambda}$ is the normalization constant. It can be easily checked that $\psi_{\lambda}(-k)=\psi_{\lambda}^{*}(k)$ since $E_{\lambda}(-k)=E_{\lambda}(k)$.
Following Ref. [29], the topology of the energy bands is obtained as
$E_{1}(k)+E_{2}(k)+E_{3}(k)=0$ (13)
$E_{1}^{2}(k)+E_{2}^{2}(k)+E_{3}^{2}(k)=2L^{2}(k)$ (14)
$E_{3}^{2}(k)-E_{1}(k)E_{2}(k)=L^{2}(k)$ (15)
$2E_{1,2}(k)=-E_{3}(k)\pm\sqrt{4L^{2}(k)-3E_{3}^{2}(k)}.$ (16)
for any given $k$. In three-dimensional Euclidean space with the axes $E_{1},E_{2},E_{3}$, Eq. (13) describes a plane passing through the origin, for which the normal line has direction cosines: $(1/\sqrt{3})(1,1,1)$. Similarly, Eq. (14) is a sphere of radius $\sqrt{2}L(k)$ centered at the origin.
Here we consider the simplest case when all the NNN hopping terms are neglected. The parameters $p$ and $q$ are simplified as $p=-(\Delta^{2}+u^{2}+v^{2}+w^{2})$ and $q=-[\Delta(u^{2}-v^{2})+uvw\cos(k)]$. The corresponding eigenvectors in this case are
$\psi_{\lambda}(k)=N_{\lambda}\left(\begin{array}[]{c}u(E_{\lambda}+\Delta)+vwe^{-ik}\\ E_{\lambda}^{2}-\Delta^{2}-w^{2}\\ v(E_{\lambda}-\Delta)+uwe^{ik}\end{array}\right).$ (17)
II.2 Symmetry properties of the Hamiltonian $\mathcal{H}(k)$
Here we will discuss three symmetry properties of the Bloch Hamiltonian $\mathcal{H}(k)$.
Time-reversal symmetry: Note that the three complex Gell-Mann matrices are odd under time reversal: $\mathcal{C}^{-1}\Lambda_{2,5,7}\mathcal{C}=-\Lambda_{2,5,7}$ and the remaining matrices are even under time reversal, with $\mathcal{C}$ being the complex conjugation operator. At the same time, the three components of the vector ${\bm{d}}$ are odd under time-reversal: $d_{2,5,7}(-k)=-d_{2,5,7}(k)$ and the remaining components are even under time reversal. Thus the Hamiltonian $\mathcal{H}(k)$ is time-reversal invariant: $\mathcal{C}\mathcal{H}(-k)\mathcal{C}=\mathcal{H}^{*}(-k)=\mathcal{H}(k)$.
[…]
Chiral symmetry: The chiral symmetry states that if there is a state $\psi_{k}$ with energy $E$ for the given Hamiltonian $H$, there must be another state $\psi_{k}^{\prime}=\Gamma\psi_{k}$ with the energy $-E$. Here $\Gamma$ is the chiral operator such that $\Gamma H\Gamma^{\dagger}=-H$. Thus the chiral operator relates the positive and negative energies and explains the electron-hole symmetry in the bands.
[…]
The concept of point chiral symmetry has been introduced for $\Delta=u_{1}=v_{1}=w_{1}=0$ case in Ref. [43, 44, 45]. Mathematically, the point chiral symmetry is stated as $\Gamma_{p}\mathcal{H}(k)\Gamma^{\dagger}=-\mathcal{H}(k+\pi)$ with the point chiral operator $\Gamma_{p}=\text{diag}(1,-1,1)$. It immediately follows $E_{1}(\pi-k)=-E_{3}(k)$ and $E_{2}(\pi-k)=-E_{2}(k)$. We extend the concept of point chiral symmetry for $\Delta\neq 0,u_{1}=v_{1}=w_{1}=0$ along with $u=v$ case. We find that the point chiral symmetry is preserved even in the presence of $\Delta$ with the condition $u=v$. The eigenstates of the shifted particle-hole symmetry are connected as $\Gamma_{p}\psi_{1}(\pi-k)=\psi_{3}(k)$ and $\Gamma_{p}\psi_{2}(\pi-k)=\psi_{2}(k)$.
The NNN hopping terms break both the chiral symmetries due to the presence of the Bloch phases ($\sim e^{\pm ik}$) arising from the inter-cell and intra-cell NNN hopping terms.
III Topological phases and bulk-boundary correspondence
In 1D, no non-trivial topological phase is possible for systems having no symmetry. For systems in odd spatial dimensions with chiral and/or inversion symmetry, three bulk-topological invariants namely the winding number ($\nu$), Zak phase ($\phi_{\lambda}$), and inversion symmetry indicator ($\chi$) are proposed [27]. These topological invariants can take non-trivial values depending on the symmetries of the system and correspond to boundary modes in the open boundary condition. The winding number is defined as follows [46]
$\nu=$ $\oint_{\rm BZ}\frac{dk}{2\pi i}\partial_{k}\log{(\det{(H(k))})},$ (18)
where $H(k)$ is the Bloch Hamiltonian. The winding number can only be quantized and is $\mathbb{Z}$-valued for systems respecting chiral symmetries.
However, for systems having inversion symmetry with or without chiral symmetry, the Zak phase can be quantized and is $\mathbb{Z}_{2}$-valued ($0,\pi$ mod $2\pi$). The Zak phase ($\phi_{\lambda}$) is defined as follows
[…]
where $\psi_{\lambda}(k)$ denotes eigenstate corresponding to the band $\lambda$.
For inversion-symmetric topological insulators with or without chiral symmetry, the Zak phase has been used to distinguish topologically trivial ($\phi_{1}=0~{}\rm mod~{}2\pi)$ and non-trivial phases ($\phi_{1}=\pi~{}\rm mod~{}2\pi$). However, there is an ambiguity in establishing the bulk-boundary correspondence. In a recent study [47], a generalization of the Zak phase namely the inter-cellular Zak phase has been proposed which establishes the bulk-boundary correspondence by correctly counting the number of boundary modes below the Fermi level. The inter-cellular Zak phase is the sum of sub-lattice Zak phases [48] which is defined as follows
[…]
where $P_{\alpha}$ denotes the projection operator on the sub-lattice $\alpha$. In the next two sub-sections, we will discuss bulk-boundary correspondence and two different topological invariants namely the Zak phase and sub-lattice winding numbers.
Figure 2: Upper panel: Dispersion of the three Bloch bands $E_{\lambda}(k)$ for various hopping terms, along with the gap closing points at $k=0,\pm\pi$ in the middle figure. Lower panel: Winding of the complex function $W_{\text{b}}\equiv W_{E_{\text{b}}}=W-E_{\text{b}}$ is shown in the $W_{{\text{b}},x}$-$W_{{\text{b}},y}$ plane with $W_{{\text{b}},x}={\rm Re}[W_{\text{b}}]$ and $W_{{\text{b}},y}={\rm Im}[W_{\text{b}}]$. The symbol $\times$ denotes the band touching at the origin of the $W_{{\text{b}},x}$-$W_{{\text{b}},y}$ plane, which corresponds to $k=0,\pm\pi$ band touching points in the $E_{\lambda}$-$k$ plane. In both the panels, we have taken $u=v=1$ and $(u_{1},v_{1},w_{1})=(0,0,0)$.
Figure 3: Upper panel: Dispersion of the three Bloch bands $E_{\lambda}(k)$ for various hopping terms, along with the gap closing points at $k=\pm\pi$ in (b) and (d). Lower panel: Winding of the complex function $W_{\text{b}}$ is shown in the $W_{{\text{b}},x}$-$W_{{\text{b}},y}$ plane. In both the panels, we have taken $u=v=1$ and $(u_{1},v_{1},w_{1})=(0,0,2)$.
Figure 4: Bulk-topological invariant: (a), (b), and (c) describe the Zak phases ($\phi_{1},~{}\phi_{2},~{}\text{and}~{}\phi_{3}$ in units of $\pi$) for $\lambda=1,~{}\lambda=2,$ $\lambda=3$ bands, respectively, in the $(w,w_{1})$-plane for $u=v=1$. (d) The sum of the Zak phases $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the $(w,w_{1})$-plane for $u=v=1$.
Figure 5: Bulk-boundary correspondence: (a) Energy band structure vs inter-cell hopping amplitude $w$ in the open boundary condition. The horizontal dashed lines are the edge modes with energy $E_{\text{b}}=\pm 1$. (b) Bulk topological invariant (sum of the Zak phases for the two bands $\phi_{1}+\phi_{3}$ (in units of $\pi$) characterizes the appearance of boundary modes and hence the topological phase transition as one tunes the inter-cell hopping term $w$ for $u=v=1$, and $w_{1}=0$. (c) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=+1$ for $(w,w_{1})=(3.52,0.0)$. (d) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=-1$ for $(w,w_{1})=(3.52,0.0)$.
Figure 6: Bulk-boundary correspondence: (a) Energy band structure vs inter-cell hopping term $w$ in the open boundary condition. The horizontal dashed lines are the edge modes with energy $E_{\text{b}}=\pm 1$. (b) Bulk topological invariant (sum of the Zak phases for the two bands $\phi_{1}+\phi_{3}$ (in units of $\pi$) characterizes the appearance of boundary modes and hence the topological phase transition as one tunes the inter-cell hopping term $w$ for $u=v=1$, and $w_{1}=2$. (c) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=+1$ for $(w,w_{1})=(3.72,2.0)$. (d) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=-1$ for $(w,w_{1})=(3.72,2.0)$.
III.1 Bulk-boundary correspondence in trimer SSH-model
Recently, the bulk-boundary correspondence in the trimer SSH model [42, 43] without NNN hopping terms has been studied. The trimer SSH model lacks full chiral symmetry but has generalized chiral, isolated, and point chiral symmetries. Therefore, the chiral winding number cannot characterize the topological phases in this system. However, the Zak phase successfully establishes the bulk-boundary condition for the system parameters respecting the inversion symmetry. Moreover, for the trimer SSH model in the absence of NNN hopping term with broken inversion symmetry, another generalization of the Zak phase namely normalized sub-lattice Zak phase has been proposed which successfully establishes the bulk-boundary condition [43].
In this section, we discuss the bulk-boundary condition in the inversion-symmetric trimer SSH lattice with NNN hopping terms ($u_{1},~{}v_{1},~{}w_{1}$) (Fig. 1). In particular, the inversion symmetry enforces the hopping to satisfy the following criteria,
$\Delta=0,u=v,~{}\text{and}~{}u_{1}=v_{1}.$ (21)
For simplicity, we will consider $u=v=1$ which sets the unit of energy and $u_{1}=v_{1}=0$ in this section. The results for different sets of $u_{1}=v_{1}\neq 0$ is given in the Appendix C.
$w_{1}=0$ case: In the absence of NNN coupling ($w_{1}=0$) and for $u=v=w$, the lower band touches the middle one at $\Gamma$-point ($k=0$) and the the middle band touches the upper one at $X$-point ($k=\pi$) simultaneously, signaling a topological phase transition (See Fig 2).
$w_{1}\neq 0$ case: For finite NNN hopping ($w_{1}\neq 0$), the degeneracy of the band touching points at $\Gamma,~{}X$-points is removed. In this case, the bulk band structure evolves according to Fig. 3 as we tune the inter-cell hopping $w$ for the fixed value of NNN hopping $w_{1}=2$. At $w=u=v$ and $w<w_{1}$, the band gap between the lower and middle bands closes at the $X$-point (See Fig.  3 (b)). Similarly, at $w>u=v$ and $w=w_{1}$, the band gap between the middle and the upper bands closes at the $X$-point (See Fig.  3 (d).
We numerically, compute the Zak phase defined in Eq. 19 for the individual bands. Figures 4 (a), (b), (c), and (d) show the variation of the Zak phases $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, and $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the ($w,w_{1}$)-plane, respectively. In the absence of NNN hopping ($w_{1}=0$), our results are in full agreement with that of Ref. [42, 43]. It can be seen that any two of the three Zak phases change in a quantized manner whenever the bulk band gap closes, which signals a topological phase transition. Accordingly, in Table 1, we label all three different topological phase transitions (topological phase transitions) characterized by the locations of the gap-closing points in the BZ while tuning $w$ and $w_{1}$.
TPT System parameters Bulk band gaps ($\Delta_{\lambda\lambda^{\prime}}$) Zak phases Sub-lattice winding number
TPT-I $w>u=v,w_{1}=0$ $u_{1}=v_{1}=0$ $\Delta_{21}$($\Delta_{32}$) closes at $\Gamma$ ($X$)-point $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ $\nu^{A}=0\leftrightarrow 2$
TPT-II $w>u=v,w<w_{1}$ $u_{1}=v_{1}\neq 0$ $\Delta_{21}$ closes only at $X$-point $\phi_{1}+\phi_{3}=0\leftrightarrow~{}\pi$ $\nu^{A}=0\leftrightarrow 1$
TPT-III $w>u=v,w>w_{1}$ $u_{1}=v_{1}\neq 0$ $\Delta_{32}$ closes only at $X$-point $\phi_{1}+\phi_{3}=\pi\leftrightarrow 2\pi$ $\nu^{A}=1\leftrightarrow 2$
Table 1: Classification of topological phase transition (TPT): Three topological phase transitions for various system parameters based on the band gaps $\Delta_{21}=E_{2}-E_{1}$ and $\Delta_{32}=E_{3}-E_{2}$ closing points in the BZ along with the topological invariant quantities such as Zak phase and sub-lattice winding number
For $w>u=v$ and $w_{1}=0$, the Zak phases $\phi_{1},\phi_{3}$ are quantized to $\pi$ while $\phi_{2}$ vanishes. In the open boundary condition, there exists at each energy $E_{\text{b}}=\pm 1$, a single boundary mode localized at each end of the system [see Fig. 5 (a) ]. These results establish the bulk-boundary correspondence. In this case the localization lengths for the boundary modes at energies $E_{\text{b}}=\pm 1$ are the same and are determined by the bulk energy gaps $\Delta_{21}$ and $\Delta_{32}$. Figure 5 (a) and (b) show the variation of the band structure in the open boundary condition and $\phi_{1}+\phi_{3}$ with respect to $w$, respectively. Figures  5 (c) and (d) show the real space profile of the wave function amplitudes for a finite open chain corresponding to the boundary modes with energies $E_{\text{b}}=+1$ and $E_{\text{b}}=-1$, respectively.
In the presence of the NNN hopping $(w_{1}\neq 0)$, the quantized change in the Zak phase $\phi_{2}$ senses both the gap closings ($\Delta_{21}$ and $\Delta_{32}$ ) at $X$ points and characterizes both the topological phase transitions (See Fig. 4 (b)). However, the quantized change in the Zak phase $\phi_{1}$ ($\phi_{3}$) senses the band touching $\Delta_{21}=0$ ( $\Delta_{32}=0$) at the $X$-points (See Fig. 4 (a) and (c)). Hence, the Zak phases $\phi_{1}$ and $\phi_{3}$ separately characterize the topological phase transition-II (TPT-II) and topological phase transition-III (TPT-III), respectively.
Figures 6 (a) and (b) show the variation of the band structure in the open boundary condition and $\phi_{1}+\phi_{3}$ with respect to $w$, respectively. Figures 6 (c) and (d) show the real space profile of wave function amplitudes for a finite open chain corresponding to the boundary modes with energies $E_{\text{b}}=+1$ and $E_{\text{b}}=-1$, respectively. Accordingly, in the open boundary condition, there exist at each energy $E_{\text{b}}=\pm 1$, a single boundary mode localized at each end of the system. In this case the localization lengths for the boundary modes at energies $E_{\text{b}}=\pm 1$ are different and are determined by the bulk energy gap $\Delta_{21}$ ($\Delta_{32}$) for negative (positive) energy boundary modes.
Based on the above analysis we observe that the addition of Zak phases $\phi_{1}+\phi_{3}$ characterizes the topological phase transitions and number of boundary modes at each end of the system. A similar analysis holds for $u_{1}=v_{1}\neq 0$. However, the phase diagram looks more complex, which is given in Appendix C (See Fig. C.7).
III.2 Sub-lattice winding numbers as bulk-topological invariants
Here we propose a different bulk topological invariant namely the sub-lattice winding number, similar to the sub-lattice Zak phases defined in Eq. 20, to discuss the bulk-boundary correspondence in this system. We construct the doubled Hermitian Hamiltonian from the chiral asymmetric Bloch Hamiltonian $H(k)$ in the following manner [49, 50]
$\tilde{H}(k)=\begin{pmatrix}0&\mathcal{H}(k)-E_{\rm b}\\ \mathcal{H}^{\dagger}(k)-E_{\rm b}&0\end{pmatrix},$ (24)
which by construction is chiral symmetric with respect to the following chiral symmetry operator
$\Sigma_{z}=\begin{pmatrix}I_{3}&0\\ 0&-I_{3}\end{pmatrix},$ (27)
where $I_{3}$ denotes the $3\times 3$ identity matrix. We choose the reference energy $E_{\rm b}$ same as the gap closing energy. The gaped nature of $\tilde{H}_{E_{\rm b}}(k)$ is defined via $\det(\tilde{H}_{E_{\rm b}}(k))\neq 0$ which is true for $\det(\mathcal{H}(k)-E_{\rm b})\neq 0$. For non-Hermitian systems, the symmetry class chiral symmetry is called class AIII, and the topological classification of insulators is given by the bulk winding numbers $\nu\in\mathbb{Z}$ which is defined as follows
$\nu_{E_{\rm b}}$ $=\oint_{\rm BZ}\frac{dk}{2\pi i}{\rm Tr}[(\mathcal{H}(k)-E_{\rm b})^{-1}\partial_{k}(\mathcal{H}(k)-E_{\rm b})].$ (28)
It is straightforward to realize that the above topological invariant vanishes for the present case. However, we can define the sub-lattice winding numbers $\nu_{E_{\rm b}}^{\alpha}$ in the following manner,
$\nu_{E_{\rm b}}^{\alpha}$ $=\oint_{\rm BZ}\frac{dk}{2\pi i}{\rm Tr}[P_{\alpha}(\mathcal{H}(k)-E_{\rm b})^{-1}\partial_{k}(\mathcal{H}(k)-E_{\rm b})P_{\alpha}],$ (29)
where $P_{\alpha}=|\alpha\rangle\langle\alpha|$ denote the projection operators for different sub-lattice sectors $\alpha=A,B,C$. Here the sub-lattice eigenstate can be expressed as $|\alpha\rangle=(\delta_{\alpha,A},\delta_{\alpha,B},\delta_{\alpha,C})^{\mathcal{T}}$ with ${\mathcal{T}}$ being the transpose operator. Although, the total winding number ($\nu_{E_{\rm b}}=\sum_{\alpha}\nu_{E_{\rm b}}^{\alpha}=0)$ vanishes for each reference energy. The sub-lattice winding numbers ($\nu_{E_{\rm b}}^{\alpha}$) can be quantized and may characterize the number of boundary modes in the system. This is indeed the case for the present system. We construct a bulk topological invariant $\nu^{\rm A}$ which is defined as follows
$\nu^{\rm A}=\sum_{E_{\rm b}}\nu_{E_{\rm b}}^{\rm A}.$ (30)
The non-trivial value of the above topological invariant $\nu^{\rm A}$ counts the number of boundary modes localized at each end of the system and establishes the bulk-boundary correspondence.
To elaborate on the above discussion, we again consider the trimer SSH chain with $u_{1}=v_{1}=0,~{}u=v=1$. In this case, the bulk energy band gap closes at ±$\pm 1$, therefore we choose the reference energies $E_{\rm b}=\pm 1$. The sub-lattice winding number $\nu_{\pm 1}^{\rm A}$ take the following form
$\nu_{\pm 1}^{\rm A}$ $=\int_{-\pi}^{\pi}\partial_{k}\log(W^{*}\pm 1)\frac{dk}{2\pi i}$
=$=\begin{cases}0,~{}&|w_{1}\pm 1|>w\\ 1,~{}&|w_{1}\pm 1|<w,\end{cases}$ (31)
where $W=w_{1}+we^{-ik}$ is nothing but the off-diagonal element of $\mathcal{H}(k)$. The sub-lattice winding number $\nu_{\pm 1}^{\rm B}$ vanishes and $\nu_{\pm 1}^{\rm C}=-\nu_{\pm 1}^{\rm A}$ due to $\sum_{\alpha}\nu_{\pm 1}^{\alpha}=0$. It is straightforward to verify from Fig. 5 and Fig. 6 that $\nu^{\rm A}=\nu_{+1}^{\rm A}+\nu_{-1}^{\rm A}$ correctly counts the number of boundary modes at each end of the system. For example, in the case of $u_{1}=v_{1}=w_{1}=0$, for $w>u=v$ and $\nu^{\rm A}=2$ determines one boundary mode with energy $E_{\rm b}=-1$ and one boundary mode with energy $E_{\rm b}=+1$ at each end of the system. In the case of $u_{1}=v_{1}$, for $w>u=v,~{}w<w_{1}$, $\nu^{\rm A}=1$ determines the one boundary mode with energy $E_{\rm b}=-1$ at each end of the system and for $w>u=v,~{}w>w_{1}$, $\nu^{\rm A}=2$ determines one boundary mode with energy $E_{\rm b}=-1$ and one boundary mode with energy $E_{\rm b}=+1$ at each end of the system.
The main advantage of this topological invariant is its analytical form and simple understanding of the topological phase transition, similar to the dimer SSH model [19]. When $k$ varies across the first BZ, the endpoint of the vector ${\bf W}_{\text{b}}$ traces out a circle of radius $w$ centered at $(w_{1}-E_{\text{b}})$ in the $W_{{\text{b}},x}$-$W_{{\text{b}},y}$ plane, as shown in the lower panels of Figs. 2 and 3. The gap closes at ${\bf W}_{\text{b}}=0$, which corresponds to the gap closing points $k=0,\pm\pi$ in $k$ space. The sub-lattice winding number will be one when the trajectory of ${\bf W}_{\text{b}}$ encloses the origin, otherwise zero. It is clear from Fig. 2 (d), (e), and (f) that the winding of the complex function $W_{\text{b}}$ is zero, ill-defined, and two for the trivial, critical, and non-trivial phases. The topological phase transition takes place at $w=u=v$ where the contour of $W_{\rm b}$ touches the origin which corresponds to the bulk-gap closing points at $k=0,\pi$. Figure 2 displays that tuning the parameter $w$ while fixing $u=v=1$ and $(u_{1},v_{1},w_{1})=(0,0,0)$ induces a topological phase transition characterized by the sub-lattice winding number from $\nu^{A}=0\leftrightarrow 2$.
Similarly, it is clear from Fig. 3 (f)-(j) that the winding of $W_{\text{b}}$ is zero, ill-defined, one, one, and two, respectively. The topological phase transitions take place at $w=u=v,w<w_{1}$ where contour of $W+1$ touches the origin which corresponds to the bulk-gap closing at $k=\pm\pi$ and also at $w=u=v,w=w_{1}$ where the contour of $W-1$ touches the origin which corresponds to the bulk-gap closing at $k=\pm\pi$. Figure 3 displays a series of topological phase transitions characterized by the sub-lattice winding number from $\nu^{A}=0\leftrightarrow 1\leftrightarrow 2$, while tuning the parameter $w$ for fixed $u=v=1$ and $(u_{1},v_{1},w_{1})=(0,0,2)$.
IV Summary and Conclusion
In this work, it is shown that the Bloch Hamiltonian of the trimer SSH model with the on-site potential energy and up to NNN hopping terms can be expressed in terms of all eight generators of the SU(3) group. The exact analytical expressions of the three dispersive energy bands and the associated eigenstates for arbitrary choices of the system parameters are provided. We discussed inversion symmetry, time-reversal symmetry, and some special chiral symmetry of the Hamiltonian. The system lacks full chiral symmetry, however, we find three different kinds of topological phase transitions while tuning the NN and NNN hopping terms. The topological phase transitions are characterized by the two bulk topological invariants, namely the Zak phase and the sub-lattice winding number. We established the bulk-boundary condition by computing the Zak phases ($\phi_{\lambda})$ for all the bands along with the boundary modes in the open boundary condition. The quantized change in two out of three Zak phases $\phi_{1}+\phi_{3}$ changing from $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ as well as $\phi_{1}+\phi_{3}=0\leftrightarrow\pi\leftrightarrow 2\pi$ characterize these topological phase transitions. We proposed a new topological invariant, the sub-lattice winding number which also characterizes the topological phase transitions. The exact analytical expressions of the sub-lattice winding number revealed that $\nu^{\alpha}$ changes from $\nu^{\alpha}=0\leftrightarrow 2$ as well as $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ across the topological phase transitions. The simple analytical method of calculating the sub-lattice winding number may help in understanding the topological phases of systems without chiral symmetry.
ACKNOWLEDGEMENT
We would like to thank Bashab Dey and SK Firoz Islam for the useful discussion. S.V. acknowledges financial support from the Institute for Basic Science in the Republic of Korea through the project IBS-R024-D1.
Appendix A Gell-Mann matrices and their basic properties
In this Appendix, we will provide well-known Gell-Mann matrices [28] and some of their basic properties. The Gell-Mann matrices are
$\Lambda_{1}$ $=$ $\left(\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ 0&0&0\end{array}\right),\;\Lambda_{2}=\left(\begin{array}[]{ccc}0&-i&0\\ i&0&0\\ 0&0&0\end{array}\right),$ (38)
$\Lambda_{3}$ $=$ $\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&0\end{array}\right),\;\Lambda_{4}=\left(\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 1&0&0\end{array}\right),$ (45)
$\Lambda_{5}$ $=$ $\left(\begin{array}[]{ccc}0&0&-i\\ 0&0&0\\ i&0&0\end{array}\right),\;\Lambda_{6}=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&1\\ 0&1&0\end{array}\right),$ (52)
$\Lambda_{7}$ $=$ $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&-i\\ 0&i&0\end{array}\right),\Lambda_{8}=\frac{1}{\sqrt{3}}\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&-2\end{array}\right).$ (59)
These matrices satisfy the Hilbert-Schmidt orthonormality condition ${\rm Tr}(\Lambda_{\mu}^{\dagger}\Lambda_{\nu})=2\delta_{\mu\nu}$.
The particles in the chain are spinless, and the time-reversal operator in the sublattice space is simply the complex conjugation operator $\mathcal{C}$. The three complex Gell-Mann matrices are odd under time reversal: $\mathcal{C}^{-1}\Lambda_{2,5,7}\mathcal{C}=-\Lambda_{2,5,7}$ and the remaining matrices are even under time reversal.
Appendix B Canonical Bloch Hamiltonian
In this Appendix, we will discuss the canonical Bloch Hamiltonian, equivalent to the periodic Bloch Hamiltonian. The canonical Bloch Hamiltonian can be obtained by using the following Fourier transformation
$(a_{l},b_{l},c_{l})$ $=$ $\frac{1}{\sqrt{N}}\sum_{k}e^{ilk}(a_{k},e^{ik/3}b_{k},e^{i2k/3}c_{k}).$
Here $a_{l}$, $b_{l}$ and $c_{l}$ are the real space field operators at the sub-lattice sites $A,B$ and $C$ in the $l$-th cell. Note that here exact locations of the sublattices within a cell $l$ are used while taking the Fourier transforms of the field operators.
The real space Hamiltonian $H$ reduces to the form
$H$ $=$ $\sum_{k}\psi_{k}^{\dagger}H(k)\psi_{k},$ (60)
where the spinor $\psi_{k}^{\dagger}=(a_{k}^{\dagger},b_{k}^{\dagger},c_{k}^{\dagger})$ and the canonical Bloch Hamiltonian $H(k)$ is given by
$H(k)$ $=$ $\left(\begin{array}[]{ccc}\Delta&e^{ik/3}U&e^{2ik/3}W\\ e^{-ik/3}U^{*}&0&e^{ik/3}V\\ e^{-2ik/3}W^{*}&e^{-ik\/3}V^{*}&-\Delta\end{array}\right).$ (64)
This Hamiltonian is periodic in $6\pi$: $H(k+6\pi)=H(k)$, due to the fact that the distance between the two successive sites within the unit cell is one-third of the unit cell size. Note that the canonical Bloch Hamiltonian $H(k)$ can be obtained from the periodic Bloch Hamiltonian $\mathcal{H}(k)$ through the following unitary transformation: $H(k)=U\mathcal{H}(k)U^{\dagger}$, where the unitary matrix $U$ is given by $U=\text{diag}(1,e^{-ik/3},e^{-i2k/3})$.
The canonical Bloch Hamiltonian $H(k)$ can be expressed in terms of the Gell-Mann matrices as
$H(k)$ $=$ ${\bm{\Lambda}}\cdot{\bm{h}(k)},$ (65)
where the components of the vector ${\bm{h}(k)}$ are $h_{1}=u\cos(k/3)+u_{1}\cos(2k/3),h_{2}=u_{1}\sin(2k/3)-u\sin(k/3),h_{3}=\Delta/2,h_{4}=w\cos(k/3)+w_{1}\cos(2k/3),h_{5}=w\sin(k/3)-w_{1}\sin(2k/3),h_{6}=v\cos(k/3)+v_{1}\cos(2k/3),h_{7}=v_{1}\sin(2k/3)-v\sin(k/3)$, and $h_{8}=\sqrt{3}\Delta/2$. The components $h_{2},h_{5},h_{7}$ are odd under time-reversal: $h_{2,5,7}(-k)=-h_{2,5,7}(k)$ and the remaining components are even under time-reversal.
Appendix C Different topological phases for $u_{1}=v_{1}\neq 0$ case
In Sec. III, we have presented results of different topological phases for $u_{1}=v_{1}=0$ case. In this Appendix, we present results of different topological phases characterized by the Zak phases ($\phi_{1}+\phi_{3}$) on $w$-$w_{1}$ parameter space for different sets of $u_{1}=v_{1}\neq 0$ case. The phase diagrams given in Fig. C.7 look more complex and reveal that the topologically non-trivial regimes corresponding to $\phi_{1}+\phi_{3}=2\pi(\pi)$ decreases (increases) with the increase of the values of $u_{1}=v_{1}$.
Figure C.7: Bulk-topological invariant: Here (a), (b), (c), (d), (e), and (f) describe the sum of the Zak phases $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the $(w,w_{1})$ plane for $(u_{1},v_{1}):(0.2,0.2),(0.5,0.5),(1.0,1.0),(1.2,1.2),(1.5,1.5)$, and $(2.0,2.0)$, respectively. In all the plots here we have taken $u=v=1$
Figure C.7: Bulk-topological invariant: Here (a), (b), (c), (d), (e), and (f) describe the sum of the Zak phases $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the $(w,w_{1})$ plane for $(u_{1},v_{1}):(0.2,0.2),(0.5,0.5),(1.0,1.0),(1.2,1.2),(1.5,1.5)$, and $(2.0,2.0)$, respectively. In all the plots here we have taken $u=v=1$
Appendix D Sub-lattice winding number for the dimer SSH model
In this Appendix, we will present the results of the sub-lattice winding number for the dimer SSH model. The Bloch Hamiltonian of the dimer SSH model is given by
$h(k)$ $=\begin{pmatrix}0&d_{k}\\ d_{k}^{*}&0\end{pmatrix},$ (68)
where $k$ denotes the Bloch wave vector and $d_{k}=t_{1}+t_{2}e^{-ik}$, with $t_{1}$ and $t_{2}$ represent nearest-neighbour intra-cell and inter-cell hopping parameters, respectively. This Hamiltonian respects time-reversal, particle-hole, chiral, and inversion symmetries. It supports edge state with energy $E_{\text{b}}=0$ at each end of the boundary.
It is straightforward to get the sub-lattice winding numbers for the sub-lattices $A$ and $B$ as
[…]
and
[…]
The sub-lattice winding numbers $\nu^{A/B}$ are integer-valued and determine the boundary modes at each end of the system. Thus the sub-lattice winding numbers $\nu^{A/B}$ successfully specify the bulk-boundary correspondence in the dimer SSH model as well.